OpenStudy (anonymous):

One month, a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month. : x = 0.6x + 384 How many people who downloaded songs from the site that month downloaded Q Sam's latest single? A.576 B. 240 C.144 D.960

OpenStudy (anonymous):

@catlover5925

OpenStudy (anonymous):

you will need to solve for x you can do these by that collecting the like terms

OpenStudy (anonymous):

if u move x to the LHS then it will be x-0.6x=384 note x is also 1x

OpenStudy (anonymous):

1x-0.6x=384 can you solve that now @Lexi_Smith11_06_14_

OpenStudy (catlover5925):

ok you have 0.6x + 384 = X the =X doesn't make sense it has to equal something in order to solve for X

OpenStudy (anonymous):